# MoFO: Momentum-Filtered Optimizer for Mitigating Forgetting in LLM Fine-Tuning

### v3

@@ B.1 Supplementary Analysis on the Top-$\alpha\%$ Filter
In this section, we provide supplementary analysis on our top-$\alpha\%$ filter, which serves as a preliminary for proving Theorem 1 in Appendix B.2.
@@
For any $z\in\mathbb{R}^{d}$, we define the top-$\alpha\%$ filter of $z$ as
@@
$S_{k}=\{i\in[d_{k}]:|z^{(k)}_{i}|\text{ ranks within the top-}\alpha\%\text{ of all }|z^{(k)}|\text{'s entries }(|z^{(k)}_{1}|,|z^{(k)}_{2}|,\dots,|z^{(k)}_{d_{k}}|)\}$
@@
To ensure that the top-$\alpha\%$ filter $\texttt{FLT}_{\alpha}(z)$ is well-defined, when multiple entries share identical absolute values and including all of them in the set $S_{k}$ would result in exceeding the $\alpha\%$ threshold of set size, the construction of $S_{k}$ prioritizes the entries with the smallest indices among those with the same absolute values.
@@
For any $z\in\mathbb{R}^{d}$, we define the $L_{1,\text{top-}\alpha\%}$ norm of $z$ as
$\|{z}\|_{1,\text{top-}\alpha\%}:=\|z\odot\texttt{FLT}_{\alpha}(z)\|_{1}.$
@@
$\|{\cdot}\|_{1,\text{top-}\alpha\%}$ is indeed a norm in ${\mathbb{R}}^{d}$.
@@
$\|{z}\|_{1,\text{top-}\alpha\%}=\|z\odot\texttt{FLT}_{\alpha}(z)\|_{1}=\sum_{k=1}^{B}\|z^{(k)}\odot{\mathbf{e}}_{S_{k}}^{(k)}\|_{1}.$ (4)
First, if $\|{z}\|_{1,\text{top-}\alpha\%}=0$, then by (4), $\|z^{(k)}\odot{\mathbf{e}}_{S_{k}}^{(k)}\|_{1}=0$ for any $1\leq k\leq B$. Thus,
@@
$\|{cz}\|_{1,\text{top-}\alpha\%}=\|cz\odot\texttt{FLT}_{\alpha}(cz)\|_{1}=c\|z\odot\texttt{FLT}_{\alpha}(z)\|_{1}=c\|{z}\|_{1,\text{top-}\alpha\%}.$
Third, for any $x,y\in\mathbb{R}^{d}$, suppose that
@@
$\|x\odot\texttt{FLT}_{\alpha}(x+y)\|_{1}=\sum_{k=1}^{B}\|x^{(k)}\odot{\mathbf{e}}^{(k)}_{S^{\prime\prime}_{k}}\leq\sum_{k=1}^{B}\|x^{(k)}\odot{\mathbf{e}}^{(k)}_{S^{\prime}_{k}}=\|x\odot\texttt{FLT}_{\alpha}(x)\|_{1}.$
@@
$\|{x+y}\|_{1,\text{top-}\alpha\%}$ $=\|(x+y)\odot\texttt{FLT}_{\alpha}(x+y)\|_{1}$
@@
$=\|{x}\|_{1,\text{top-}\alpha\%}+\|{y}\|_{1,\text{top-}\alpha\%}.$
@@
By Proposition 1, $\|{\cdot}\|_{1,\text{top-}\alpha\%}$ is a norm in ${\mathbb{R}}^{d}$, so we have
@@
$=\|{x}\|_{1,\text{top-}\alpha\%}-\|{y}\|_{1,\text{top-}\alpha\%}+\|y\odot\texttt{FLT}_{\alpha}(y)\|_{1}-\|x\odot\texttt{FLT}_{\alpha}(y)\|_{1}$
$\leq\|{x-y}\|_{1,\text{top-}\alpha\%}+\|(y-x)\odot\texttt{FLT}_{\alpha}(y)\|_{1}$
@@
B.2 Proof of Theorem 1
Our proof of Theorem 1 follows the convergence analysis of the full-batch Adam optimizer in (Shi et al., 2021), with novel adaptations to address the unique aspects of MoFO.
@@ B.2 Proof of Theorem 1
By the update rules of the first and second moment estimates
@@
By mathematical induction, for any $1\leq i\leq d$, we have
@@
where $C=\frac{\sqrt{d\cdot(\alpha\%)+B}}{\sqrt{1-\beta_{2}}(1-\beta_{1}/\sqrt{\beta_{2}})}$.
@@
By Equation (5) and (6), we get
@@
MoFO actually choose $\lceil d_{k}\times\alpha\%\rceil$ entries to update in each part $k$ of parameters. Then for any $z\in{\mathbb{R}}^{d}$, we have
#$\#\{1\leq i\leq d:\texttt{FLT}_{\alpha}(z)_{i}=1\}=\sum_{k=1}^{B}\lceil d_{k}\cdot(\alpha\%)\rceil\leq\sum_{k=1}^{B}(d_{k}\cdot(\alpha\%)+1)=d\cdot(\alpha\%)+B.$
Then for the $L_{2}$-distance, we have
@@
$\leq\frac{\sqrt{d\cdot(\alpha\%)+B}}{\sqrt{1-\beta_{2}}(1-\beta_{1}/\sqrt{\beta_{2}})}\cdot\eta_{t}$
@@
Since $\nabla\mathcal{L}$ has Lipschitz constant $L$, we get
$|g_{i,t}-g_{i,s}|\leq\|g_{t}-g_{s}\|_{2}=\|\nabla\mathcal{L}(\theta_{t-1})-\nabla\mathcal{L}(\theta_{t-1})\|_{2}\leq L\|\theta_{t-1}-\theta_{s-1}\|_{2}.$ (7)
By Lemma 2, for any $t>s\geq 1$, we have
@@
Then we have
@@
For the second moment estimate, we have
@@
Therefore,
@@
$\left\|\frac{m_{t}}{1-\beta_{1}^{t}}-g_{t}\right\|_{1}\leq\frac{2\sqrt{2}\beta_{1}\sqrt{d}LC\eta}{(1-\beta_{1})^{2}\sqrt{t}}.$
@@
$\left\|\frac{m_{t}}{1-\beta_{1}^{t}}-g_{t}\right\|_{2}$ $\leq\frac{1-\beta_{1}}{1-\beta_{1}^{t}}\sum_{s=1}^{t}\beta_{1}^{t-s}\|g_{t}-g_{s}\|_{2}\leq\sum_{s=1}^{t}\beta_{1}^{t-s}\|g_{t}-g_{s}\|_{2}$
@@
$\left\|\frac{m_{t}}{1-\beta_{1}^{t}}-g_{t}\right\|_{1}\leq\sqrt{d}\left\|\frac{m_{t}}{1-\beta_{1}^{t}}-g_{t}\right\|_{2}\leq\frac{2\sqrt{2}\beta_{1}\sqrt{d}LC\eta}{(1-\beta_{1})^{2}\sqrt{t}}.$
@@
$\leq-\frac{\sqrt{1-\beta_{2}}\cdot\eta}{\sqrt{t}}\|g_{t}\odot\texttt{FLT}_{\alpha}(m_{t})\|_{1}+\left[\frac{2\sqrt{2}\beta_{1}\sqrt{1-\beta_{2}}}{(1-\beta_{1})^{2}}+\frac{4}{\sqrt{1-\beta_{2}}}+\frac{C}{2}\right]\frac{LC\eta^{2}(d\cdot(\alpha\%)+B)}{t}.$
@@
$\leq 2\left\|g_{t}-\frac{m_{t}}{1-\beta_{1}^{t}}\right\|_{1}$
@@
Then we get
$\|g_{t}\|_{\infty}$ $\leq\|{g_{t}}\|_{1,\text{top-}\alpha\%}=\|g_{t}\odot\texttt{FLT}_{\alpha}(g_{t})\|_{1}$ (14)
$\leq\|g_{t}\odot\texttt{FLT}_{\alpha}(m_{t})\|_{1}+\frac{4\sqrt{2}\beta_{1}\sqrt{d}LC\eta}{(1-\beta_{2})^{2}\sqrt{t}}.$
@@
$\leq-\frac{\sqrt{1-\beta_{2}}\cdot\eta}{\sqrt{t}}\|g_{t}\|_{\infty}+\left[\frac{2\sqrt{2}\beta_{1}\sqrt{1-\beta_{2}}}{(1-\beta_{1})^{2}}+\frac{4}{\sqrt{1-\beta_{2}}}+\frac{C}{2}\right]\frac{LC\eta^{2}(d\cdot(\alpha\%)+B)}{t}+\frac{4\sqrt{2}\beta_{1}\sqrt{d}LC\eta^{2}}{(1-\beta_{2})^{\frac{3}{2}}t}$
$=-\frac{C_{1}}{\sqrt{t}}\|g_{t}\|_{\infty}+\frac{C_{2}}{t}$
$\leq-\frac{C_{1}}{\sqrt{t}}\min_{1\leq t\leq T}\|g_{t}\|_{\infty}+\frac{C_{2}}{t},$
@@
$C_{2}$ $=LC\eta^{2}\cdot\left\{\left[\frac{2\sqrt{2}\beta_{1}\sqrt{1-\beta_{2}}}{(1-\beta_{1})^{2}}+\frac{4}{\sqrt{1-\beta_{2}}}+\frac{C}{2}\right](d\cdot(\alpha\%)+B)+\frac{4\sqrt{2}\beta_{1}\sqrt{d}}{(1-\beta_{2})^{\frac{3}{2}}}\right\}.$
@@
$\leq-C_{1}\left(\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\right)\cdot\min_{1\leq t\leq T}\|g_{t}\|_{\infty}+C_{2}\sum_{t=1}^{T}\frac{1}{t}.$
@@
$\min_{0\leq t\leq T-1}\|\nabla\mathcal{L}(\theta_{t})\|_{\infty}$ $=\min_{1\leq t\leq T}\|g_{t}\|_{\infty}$
$\leq\frac{\mathcal{L}(\theta_{0})-\mathcal{L}^{*}+C_{2}\sum_{t=1}^{T}\frac{1}{t}}{C_{1}\sum_{t=1}^{T}\frac{1}{\sqrt{t}}}\leq\frac{\mathcal{L}(\theta_{0})-\mathcal{L}^{*}+C_{2}(1+\log T)}{2C_{1}(\sqrt{T+1}-1)}$
$=\mathcal{O}\left(\frac{\log T}{\sqrt{T}}\right).$
B.3 Proof of Theorem 2

### v4

@@ A.1 Supplementary Analysis on the Top-$\alpha$ Filter
In this section, we provide supplementary analysis on our top-$\alpha$ filter, which serves as a preliminary for proving Theorem 1 in Appendix A.2.
@@
For any $z\in\mathbb{R}^{d}$, we define the top-$\alpha$ filter of $z$ as
@@
[…]
@@
To ensure that the top-$\alpha$ filter $\texttt{FLT}_{\alpha}(z)$ is well-defined, when multiple entries share identical absolute values and including all of them in the set $S_{k}$ would result in exceeding the $\alpha$ threshold of set size, the construction of $S_{k}$ prioritizes the entries with the smallest indices among those with the same absolute values.
@@
For any $z\in\mathbb{R}^{d}$, we define the $L_{1,\text{top-}\alpha}$ norm of $z$ as
$\|{z}\|_{1,\text{top-}\alpha}:=\|z\odot\texttt{FLT}_{\alpha}(z)\|_{1}.$
@@
$\|{\cdot}\|_{1,\text{top-}\alpha}$ is indeed a norm in ${\mathbb{R}}^{d}$.
@@
$\|{z}\|_{1,\text{top-}\alpha}=\|z\odot\texttt{FLT}_{\alpha}(z)\|_{1}=\sum_{k=1}^{B}\|z^{(k)}\odot{\mathbf{e}}_{S_{k}}^{(k)}\|_{1}.$ (6)
First, if $\|{z}\|_{1,\text{top-}\alpha}=0$, then by (6), $\|z^{(k)}\odot{\mathbf{e}}_{S_{k}}^{(k)}\|_{1}=0$ for any $1\leq k\leq B$. Thus,
@@
$\|{cz}\|_{1,\text{top-}\alpha}=\|cz\odot\texttt{FLT}_{\alpha}(cz)\|_{1}=c\|z\odot\texttt{FLT}_{\alpha}(z)\|_{1}=c\|{z}\|_{1,\text{top-}\alpha}.$
Third, for any $x,y\in\mathbb{R}^{d}$, we let
[…]
$S_{k}^{\prime\prime}$ $=\{i\in[d_{k}]:|x^{(k)}_{i}+y^{(k)}_{i}|\text{ ranks within the top-}\alpha\text{ of all }$
[…]
Then we have
@@
$\|x\odot\texttt{FLT}_{\alpha}(x+y)\|_{1}=\sum_{k=1}^{B}\|x^{(k)}\odot{\mathbf{e}}^{(k)}_{S^{\prime\prime}_{k}}\|_{1}\leq\sum_{k=1}^{B}\|x^{(k)}\odot{\mathbf{e}}^{(k)}_{S^{\prime}_{k}}\|_{1}=\|x\odot\texttt{FLT}_{\alpha}(x)\|_{1}.$
@@
$\|{x+y}\|_{1,\text{top-}\alpha}$ $=\|(x+y)\odot\texttt{FLT}_{\alpha}(x+y)\|_{1}$
@@
$=\|{x}\|_{1,\text{top-}\alpha}+\|{y}\|_{1,\text{top-}\alpha}.$
@@
By Proposition 1, $\|{\cdot}\|_{1,\text{top-}\alpha}$ is a norm in ${\mathbb{R}}^{d}$, so we have
@@
$=\|{x}\|_{1,\text{top-}\alpha}-\|{y}\|_{1,\text{top-}\alpha}+\|y\odot\texttt{FLT}_{\alpha}(y)\|_{1}-\|x\odot\texttt{FLT}_{\alpha}(y)\|_{1}$
$\leq\|{x-y}\|_{1,\text{top-}\alpha}+\|(y-x)\odot\texttt{FLT}_{\alpha}(y)\|_{1}$
@@
The lemma below quantifies the relationship between $L_{1,\text{top-}\alpha}$ and $L_{p}$ norms for $p\in[1,+\infty]$.
Lemma 2.
Assume the parameter space ${\mathbb{R}}^{d}$ is decomposed into $B$ blocks ${\mathbb{R}}^{d_{1}}\times\cdots\times{\mathbb{R}}^{d_{B}}$ and $z=\mathrm{Concat}(z^{(1)},\dots,z^{(B)})$ with $z^{(k)}\in{\mathbb{R}}^{d_{k}}$. Then for any $z\in{\mathbb{R}}^{d}$ and any $p\in[1,\infty]$, it holds that
$\alpha\|z\|_{p}\leq\|{z}\|_{1,\text{top-}\alpha}\leq(d\alpha+B)^{1-\frac{1}{p}}\|z\|_{p}.$
Proof.
We first prove the $p=1$ case. Fix any block $k$. Write the absolute values in nonincreasing order $a_{1}^{(k)}\geq\cdots\geq a_{d_{k}}^{(k)}\geq 0$, where $a_{i}^{(k)}:=|z_{i}^{(k)}|$. Let $m_{k}:=|S_{k}(z)|$; by the definition of the top-$\alpha$ filter (Definition 1) and the tie-breaking rule in Remark 2, $m_{k}=\lceil\alpha d_{k}\rceil$. Since the average of the top $m_{k}$ numbers is at least the overall average, we have
$\sum_{i\in S_{k}(z)}|z_{i}^{(k)}|\geq\frac{m_{k}}{d_{k}}\sum_{i=1}^{d_{k}}|z_{i}^{(k)}|\geq\alpha\|z^{(k)}\|_{1}.$
Summing over $k$ yields $\|{z}\|_{1,\text{top-}\alpha}\geq\alpha\sum_{k}\|z^{(k)}\|_{1}=\alpha\|z\|_{1}$. The upper bound $\|{z}\|_{1,\text{top-}\alpha}\leq\|z\|_{1}$ is immediate since $\texttt{FLT}_{\alpha}(z)$ is a $\{0,1\}$ filter.
Next, let $p\in(1,+\infty]$. The lower bound follows from
$\|{z}\|_{1,\text{top-}\alpha}\geq\alpha\|z\|_{1}\geq\alpha\|z\|_{p},$
since $\|z\|_{1}\geq\|z\|_{p}$. For the upper bound, fix $k$ and apply Hölder’s inequality on the subset $S_{k}(z)$:
$\sum_{i\in S_{k}(z)}|z_{i}^{(k)}|\leq\Big(\sum_{i\in S_{k}(z)}1^{q}\Big)^{1/q}\Big(\sum_{i=1}^{d_{k}}|z_{i}^{(k)}|^{p}\Big)^{1/p}=m_{k}^{1/q}\|z^{(k)}\|_{p}=m_{k}^{1-\frac{1}{p}}\|z^{(k)}\|_{p},$
where $q=\frac{p}{p-1}$ and we used 1/q=11/p1/q=1-1/p. Summing over $k$ and applying Hölder again to the finite sum $\sum_{k=1}^{B}m_{k}^{1-1/p}\|z^{(k)}\|_{p}$ gives
$\|{z}\|_{1,\text{top-}\alpha}=\sum_{k=1}^{B}\sum_{i\in S_{k}(z)}|z_{i}^{(k)}|\leq\sum_{k=1}^{B}m_{k}^{1-\frac{1}{p}}\|z^{(k)}\|_{p}\leq\Big(\sum_{k=1}^{B}m_{k}\Big)^{1-\frac{1}{p}}\Big(\sum_{k=1}^{B}\|z^{(k)}\|_{p}^{p}\Big)^{1/p}.$
We set $m:=\sum_{k=1}^{B}m_{k}$. Since $\sum_{k=1}^{B}\|z^{(k)}\|_{p}^{p}=\|z\|_{p}^{p}$ (blocks are disjoint) and $m=\sum_{k}m_{k}=\sum_{k}\lceil\alpha d_{k}\rceil\leq\alpha d+B$, we obtain
$\|{z}\|_{1,\text{top-}\alpha}\leq m^{1-\frac{1}{p}}\|z\|_{p}\leq(\alpha d+B)^{1-\frac{1}{p}}\|z\|_{p}.$
For $p=+\infty$ the same argument applies with $q=1$ and yields $\sum_{i\in S_{k}(z)}|z_{i}^{(k)}|\leq m_{k}\|z^{(k)}\|_{\infty}$, hence $\|{z}\|_{1,\text{top-}\alpha}\leq m\|z\|_{\infty}$; the lower bound $\|{z}\|_{1,\text{top-}\alpha}\geq\alpha\|z\|_{\infty}$ follows from $\|z\|_{1}\geq\|z\|_{\infty}$. This completes the proof. ∎
A.2 Proof of Theorem 1 (Convergence of MoFO)
Notation recap (Appendix A.2).
  • $g_{t}=\nabla\mathcal{L}(\theta_{t-1})\in\mathbb{R}^{d}$ is the full-batch gradient at step $t$.
  • Step size: $\eta_{t}=\eta/\sqrt{t}$; hyperparameters satisfy $\beta_{1}<\sqrt{\beta_{2}}<1$.
  • First/second moments: $m_{t},v_{t}\in\mathbb{R}^{d}$ with updates
    $(m_{i,t},v_{i,t})=\big((1-\beta_{1})g_{i,t}+\beta_{1}m_{i,t-1},\,(1-\beta_{2})g_{i,t}^{2}+\beta_{2}v_{i,t-1}\big);$
    bias-corrected first/second moments: $\hat{m}_{t}=m_{t}/(1-\beta_{1}^{t})$, $\hat{v}_{t}=v_{t}/(1-\beta_{2}^{t})$.
  • MoFO’s filter $\texttt{FLT}_{\alpha}(\cdot)\in\{0,1\}^{d}$: in each partition $k\in[B]$ of size $d_{k}$ (with $\sum_{k}d_{k}=d$), we keep $\lceil d_{k}\alpha\rceil$ entries and zero out the others. We write $\|z\odot\texttt{FLT}_{\alpha}(z)\|_{1}$ for the $\ell_{1}$-norm of the kept coordinates and $\|{x}\|_{1,\text{top-}\alpha}\triangleq\|x\odot\texttt{FLT}_{\alpha}(x)\|_{1}$.
  • MoFO’s update: $\theta_{i,t}-\theta_{i,t-1}=-\,\eta_{t}\,\hat{m}_{i,t}/\sqrt{\hat{v}_{i,t}}$ if $\texttt{FLT}_{\alpha}(m_{t})_{i}=1$, and 0 otherwise.
Our proof of Theorem 1 follows the convergence analysis of the full-batch Adam optimizer in Shi et al. (2021), with novel adaptations to address the unique aspects of MoFO.
@@ A.2 Proof of Theorem 1 (Convergence of MoFO)
By the update rules of the first and second moment estimates, we have
@@
So by mathematical induction, for any $1\leq i\leq d$, we have
@@
where $C=\frac{\sqrt{d\alpha+B}}{\sqrt{1-\beta_{2}}(1-\beta_{1}/\sqrt{\beta_{2}})}$.
@@
By Equation (7) and (8), the first/second moments $m_{i,t},v_{i,t}$ satisfy
@@
The parameter vector is partitioned into $B$ blocks with sizes $\{d_{k}\}_{k=1}^{B}$ and MoFO actually choose $\lceil d_{k}\alpha\rceil$ entries to update in each part $k$ of parameters. Then for any $z\in{\mathbb{R}}^{d}$, we have
$\{1\leq i\leq d:\texttt{FLT}_{\alpha}(z)_{i}=1\}=\sum_{k=1}^{B}\lceil d_{k}\alpha\rceil\leq\sum_{k=1}^{B}(d_{k}\alpha+1)=d\alpha+B.$
Then for the $L_{2}$-norm of the parameter update, we have
@@
$\leq\frac{\sqrt{d\alpha+B}}{\sqrt{1-\beta_{2}}(1-\beta_{1}/\sqrt{\beta_{2}})}\cdot\eta_{t}$
@@
$|g_{i,t}-g_{i,s}|\leq\|g_{t}-g_{s}\|_{2}\leq\frac{2\sqrt{2}LC\eta(t-s)}{\sqrt{t}},$
where $C=\frac{\sqrt{d\alpha+B}}{\sqrt{1-\beta_{2}}(1-\beta_{1}/\sqrt{\beta_{2}})}$.
@@
Fix the iterations steps $t$ and $s$ with $t\geq s\geq 1$. Since $\nabla\mathcal{L}$ has Lipschitz constant $L$, the gradient difference between the step $t$ and $s$ satisfies
$|g_{i,t}-g_{i,s}|\leq\|g_{t}-g_{s}\|_{2}=\|\nabla\mathcal{L}(\theta_{t-1})-\nabla\mathcal{L}(\theta_{s-1})\|_{2}\leq L\|\theta_{t-1}-\theta_{s-1}\|_{2}.$ (9)
By Lemma 3, for any $t>s\geq 1$, the parameter difference satisfies
@@
Then for the product of gradient and momentum, we have
@@
For the second moment $v_{i,t}$, we have
@@
Therefore, for the bias-corrected first/second moments $\hat{m}_{i,t}$ and $\hat{v}_{i,t}$, it holds that
@@
$\left\|\hat{m}_{t}-g_{t}\right\|_{1}\leq\frac{2\sqrt{2}\beta_{1}\sqrt{d}LC\eta}{(1-\beta_{1})^{2}\sqrt{t}}.$
@@
$\|\hat{m}_{t}-g_{t}\|_{2}=\left\|\frac{m_{t}}{1-\beta_{1}^{t}}-g_{t}\right\|_{2}$ $\leq\frac{1-\beta_{1}}{1-\beta_{1}^{t}}\sum_{s=1}^{t}\beta_{1}^{t-s}\|g_{t}-g_{s}\|_{2}\leq\sum_{s=1}^{t}\beta_{1}^{t-s}\|g_{t}-g_{s}\|_{2}$
@@
$\|\hat{m}_{t}-g_{t}\|_{1}\leq\sqrt{d}\|\hat{m}_{t}-g_{t}\|_{2}\leq\frac{2\sqrt{2}\beta_{1}\sqrt{d}LC\eta}{(1-\beta_{1})^{2}\sqrt{t}}.$
@@
$\leq-\frac{\sqrt{1-\beta_{2}}\cdot\eta}{\sqrt{t}}\|g_{t}\odot\texttt{FLT}_{\alpha}(m_{t})\|_{1}+\left[\frac{2\sqrt{2}\beta_{1}\sqrt{1-\beta_{2}}}{(1-\beta_{1})^{2}}+\frac{4}{\sqrt{1-\beta_{2}}}+\frac{C}{2}\right]\frac{LC\eta^{2}(d\alpha+B)}{t}.$
@@
$=\|g_{t}\odot\texttt{FLT}_{\alpha}(\hat{m}_{t})\|_{1}$
$\leq 2\left\|g_{t}-\hat{m}_{t}\right\|_{1}$
@@
$\leq-\frac{\sqrt{1-\beta_{2}}\cdot\eta}{\sqrt{t}}\|g_{t}\odot\texttt{FLT}_{\alpha}(g_{t})\|_{1}+\left[\frac{2\sqrt{2}\beta_{1}\sqrt{1-\beta_{2}}}{(1-\beta_{1})^{2}}+\frac{4}{\sqrt{1-\beta_{2}}}+\frac{C}{2}\right]\frac{LC\eta^{2}(d\alpha+B)}{t}$
$\qquad\qquad\qquad\qquad\qquad+\frac{4\sqrt{2}\beta_{1}\sqrt{d}LC\eta^{2}}{(1-\beta_{2})^{\frac{3}{2}}t}$
$=-\frac{C_{1}}{\sqrt{t}}\|{g_{t}}\|_{1,\text{top-}\alpha}+\frac{C_{2}}{t}\leq-\frac{C_{1}}{\sqrt{t}}\min_{1\leq t\leq T}\|{g_{t}}\|_{1,\text{top-}\alpha}+\frac{C_{2}}{t},$
@@
$C_{2}$ $=LC\eta^{2}\cdot\left\{\left[\frac{2\sqrt{2}\beta_{1}\sqrt{1-\beta_{2}}}{(1-\beta_{1})^{2}}+\frac{4}{\sqrt{1-\beta_{2}}}+\frac{C}{2}\right](d\alpha+B)+\frac{4\sqrt{2}\beta_{1}\sqrt{d}}{(1-\beta_{2})^{\frac{3}{2}}}\right\}.$
@@
$\leq-C_{1}\left(\sum_{t=1}^{T}\frac{1}{\sqrt{t}}\right)\cdot\min_{1\leq t\leq T}\|g_{t}\odot\texttt{FLT}_{\alpha}(g_{t})\|_{1}+C_{2}\sum_{t=1}^{T}\frac{1}{t}.$
@@
$\min_{0\leq t\leq T-1}\|{\nabla\mathcal{L}(\theta_{t})}\|_{1,\text{top-}\alpha}$ $=\min_{1\leq t\leq T}\|{g_{t}}\|_{1,\text{top-}\alpha}=\min_{1\leq t\leq T}\|g_{t}\odot\texttt{FLT}_{\alpha}{(g_{t})}\|_{1}$
$\leq\frac{\mathcal{L}(\theta_{0})-\mathcal{L}^{*}+C_{2}\sum_{t=1}^{T}\frac{1}{t}}{C_{1}\sum_{t=1}^{T}\frac{1}{\sqrt{t}}}$
$\leq\frac{\mathcal{L}(\theta_{0})-\mathcal{L}^{*}+C_{2}(1+\log T)}{2C_{1}(\sqrt{T+1}-1)}.$
Thus, we have
$\min_{0\leq t\leq T-1}\|{\nabla\mathcal{L}(\theta_{t})}\|_{1,\text{top-}\alpha}=\min_{1\leq t\leq T}\|\mathcal{L}(\theta_{t})\odot\texttt{FLT}_{\alpha}{(\mathcal{L}(\theta_{t}))}\|_{1}=\mathcal{O}\left(\frac{\log T}{\sqrt{T}}\right).$
By the relationship between $L_{1,\text{top-}\alpha}$ norm and $L_{p}$ norm in Lemma 2, for any $p\in[1,+\infty]$,
$\min_{0\leq t\leq T-1}\|\nabla\mathcal{L}(\theta_{t})\|_{p}\leq\frac{1}{\alpha}\min_{0\leq t\leq T-1}\|{\nabla\mathcal{L}(\theta_{t})}\|_{1,\text{top-}\alpha}.$
Therefore,
$\min_{0\leq t\leq T-1}\|\nabla\mathcal{L}(\theta_{t})\|_{p}=\mathcal{O}\left(\frac{\log T}{\sqrt{T}}\right).$
A.3 Proof of Theorem 2 (Illustrative example: forgetting mitigation of MoFO)